OpenStudy (anonymous):

what is the similarity ratio of a cube with the volume 216 m^3 to a cube with volume 2,744 m^3 ? 3:7 7:3 36:196 196:36 The volumes of two similar solids are 1,728 m^3 and 343 m^3. The surface area of the larger solid is 576 m^2. What is the surface area of the smaller solid? 196 m^2 76 m^2 1,372 m^2 392 m^2

OpenStudy (ranga):

Assume smaller cube has side s and larger cube has side S. Volume of a cube is side^3. Ratio of volumes is: s^3 / S^3 = 216 / 2744 = 6^3 / 14^3 Take cube root on both sides: s / S = 6 / 14 = 3 / 7 s : S = 3 : 7

OpenStudy (anonymous):

I dont get it

OpenStudy (ranga):

They have given you the volumes of two cubes and they are asking what is the ratio of their sides. If the length of the side of a cube is x, its volume is x^3. So if you know the volume, you can find the length of the side by taking the cube root of the volume. Volume of smaller cube = 216. Length of the side of the smaller cube = cube root of 216 = 6 Volume of bigger cube = 2744. Length of the side of the bigger cube = cube root of 2744 = 14 The ratio of the sides is; 6:14 = 3:7